# Bitangent surfaces and involutions of quartic surfaces

### v2

@@ 3. Quartic surfaces: $p\neq 2$
In the special case $d=4$, we expect that, for a smooth or nodal quartic surface in characteristic zero, the bidegree $(m,n)$ of ${\mathrm{Bit}}(X)$ is equal to $(12,28)$. Let $S$ be an irreducible component of ${\mathrm{Bit}}(X)$ and let $q_{S}:Z_{S}\to S$ be the restriction of the tautological line subbundle $q:Z\to G_{1}(\mathbb{P}^{3})$ to the congruence $S$. The pre-image of $X$ under the projection $p_{S}:Z_{S}\to\mathbb{P}^{3}$ is a double cover of $S$ defined by the projection $q_{S}$:
@@
Thus, any irreducible component $S$ of ${\mathrm{Bit}}(X)$ with non-zero order and non-zero class defines a birational involution $\sigma_{S}$ of $X$. Its set of fixed points is equal to the pre-image of the locus of points on $X$ such that there exists a line intersecting $X$ at this point with multiplicity $4$, or contained in it.
@@
It follows from Kummer’s classification of congruences of order $2$ without fundamental curves that the moduli space of quartic surfaces with $16\geq\mu\geq 11$ nodes contains an irreducible component such that a general surface from this component has reducible congruence ${\mathrm{Bit}}(X)$. These irreducible components are realized as the focal surfaces of the congruences of order 2 without fundamental curves. The number of irreducible components and their bidegree are given in the following table (see Kummer [15], Sturm [23, Art 348]).
@@
The irreducible components of order 2 are confocal congruences, i.e. congruences of lines with the same focal surface.
@@
Let us explain how to see these six irreducible components of bidegree $(2,2)$. Recall that $X$ admits a smooth model $\tilde{X}$ as a surface of degree $8$ in $\mathbb{P}^{5}$ (see [3, 10.3.3]). The surface $\tilde{X}$ is a complete intersection of three quadrics in $\mathbb{P}^{5}$. One can choose projective coordinates such that $\tilde{X}$ is given by the following equations:
@@
$[z_{1},z_{2},z_{3},z_{4},z_{5},z_{6}]\to[z_{1},\pm z_{2},\pm z_{3},\pm z_{4},\pm z_{5},\pm z_{6}]$
induce automorphisms of $\tilde{X}$ which generate a $2$-elementary group of order $2^{5}$. It contains a subgroup of index $2$ that consists of the identity and $15$ projective involutions that restrict to symplectic automorphisms of $\tilde{X}$. [^1] This means that the involution leaves invariant a non-zero regular $2$-form on the surface, or, equivalently, the set of its fixed points is non-empty and finite. They are the involutions that change exactly two signs at the coordinates. The quotient surface of $\tilde{X}$ by a symplectic automorphism is a K3 surface with rational double points. The remaining sixteen involutions are decomposed into two sets. One set consists of ten involutions that change exactly three signs at the coordinates. They are fixed-point-free with the orbit space isomorphic to an Enriques surface. Another set consists of six involutions $g_{i}$ that change only one coordinate with the orbit space isomorphic to a del Pezzo surface of degree $4$. These are the involutions we are interested in.
The first quadric $V(\sum z_{i}^{2})$ can be identified with the Grassmannian quadric $\mathbb{G}$ written in Klein coordinates $z_{i}$ corresponding to six apolar line complexes (see, for example, [3, 10.2]). It is known that any automorphism of $G_{1}(\mathbb{P}^{3})$ comes from a projective collineation or correlation. The involutions $g_{i}$ preserve the apolar line complexes $V(z_{i})$, and hence, come from a correlation. They transform points to planes, and therefore, act on $\mathbb{G}$ by transforming the plane of lines through a point (an $\alpha$-plane) to the plane of lines contained in a given plane (a $\beta$-plane).
@@
The surface $\tilde{X}$ sits in $G_{1}(\mathbb{P}^{3})$ as a singular surface of the quadratic line complex $Y=V(\sum a_{i}z_{i}^{2})\cap\mathbb{G}$, that is, it parameterizes the lines that are singular points of the intersection of $Y$ with $\alpha$-planes. We can identify each such $\alpha$-plane $\Omega(x)$ with the point $x\in X$. An automorphism $g_{i}$ acts on $X$ by sending $\Omega(x)$ to $\Omega(\Pi)$ Let $x\in X$ be a general point and let $\Pi=\mathbb{T}_{x}(X)$ be the embedded tangent space of $X$ at $x$. As we explained above, $g_{i}$ interchanges two families $\{\Omega(x)\}_{x\in{\bf P}^{3}}$, $\{\Omega(\Pi)\}_{\Pi\subset\mathbb{P}^{3}}$ of planes, and hence,
@@
$\emptyset\neq\Omega(x)\cap H_{i}=g_{i}(\Omega(x)\cap H_{i})=\Omega(\Pi^{\prime})\cap H_{i}=\Omega(x)\cap\Omega(\Pi^{\prime}),$
$\emptyset\neq\Omega(\Pi)\cap H_{i}=g_{i}(\Omega(\Pi)\cap H_{i})=\Omega(\Pi^{\prime})\cap H_{i}=\Omega(x^{\prime})\cap\Omega(\Pi).$
@@
Also, note that the fixed locus of the involution $g_{i}$ is a canonical curve of genus $5$ given as the intersection of three diagonal quadrics. It has special group of automorphisms isomorphic to the $2$-elementary group $2^{4}$. It is classically known as a Humbert curve. Its image in the quartic del Pezzo surface $\mathcal{D}$ belongs to $|-2K_{\mathcal{D}}|$. It is the branch curve of the double cover $\tilde{X}\to\mathcal{D}$. It has the distinguished property that any of the 16 lines on $\mathcal{D}$ splits into a pair of lines on $\tilde{X}$.
@@ 4. Birational involutions of a quartic surface
Let $\sigma$ be a birational involution of a quartic surface $X$. The closure of lines spanned by orbits of $\sigma$ is an irreducible congruence $S(\sigma)$ of lines. Fix a general point $P\in\mathbb{P}^{3}$, then the order of $S(\sigma)$ is equal to the cardinality of the set $\{x\in\mathbb{P}^{3}:P\in\langle x,\sigma(x)\rangle\}$.
@@
where $P=[a,b,c,d],x=[x_{0},x_{1},x_{2},x_{3}],T(x)=[y_{0},y_{1},y_{2},y_{3}].$ It is expected to be a curve of degree $12k+$ 1, where $k$ is the algebraic degree of $T$ [10, p. 175]. However, it also includes the locus of fundamental points of $T$, as well as the closure of the locus of fixed points of $T$. The class of $S(\sigma)$ is expected to be equal to #$\#\Pi\cap T(\Pi)\cap X=4k$.
@@
Let $S(\sigma)\subset G_{1}(\mathbb{P}^{3})$ be the congruence of lines, the image of the rational map $\phi_{\sigma}$.
@@
Here we give an example of an involution of Type (ii). Let $\sigma$ be the restriction of a projective involution $T$ in $\mathbb{P}^{3}$. Suppose that $p\neq 2$ and the fixed locus of $T$ is the union of two skew lines $\ell_{1},\ell_{2}$. It is clear that the lines $\ell_{x}$ are invariant lines of $T$, hence each such a line has two fixed points on it, one on each of the lines $\ell_{1},\ell_{2}$. This shows that the congruence $S(\sigma)$ is of bidegree $(1,1)$ isomorphic to a smooth quadric. Since the set of fixed points $X^{\sigma}$ consists of $8$ isolated fixed points or it is one-dimensional, we obtain that either both lines lie in $X$ or each line intersects $X$ at four distinct points. In the first case, the degree of $\phi_{\sigma}$ is equal to one. Assume $X$ is smooth, then $Y=X/(\sigma)$ is also smooth. The cover $X\to Y=X/(\sigma)$ is ramified over two $(-2)$-curves, and its branch curve is the union of two disjoint smooth rational curves $C_{1},C_{2}$ with self-intersection $-4$. By the adjunction formula, $|-K_{Y}|=\emptyset,|-2K_{Y}|=\{C_{1}+C_{2}\}$. Thus, $Y$ is a Coble surface obtained by blowing up 10 points on a smooth quadric $Q$, the eight intersection points of two nodal quartic curves $\bar{C}_{1},\bar{C_{2}}$ of bidegree $(2,2)$. The pre-images of the ten exceptional curves on $Y$ in $X$ are ten invariant lines lying in $X$ that contain infinitely many orbits of $\sigma$.
@@
We check that the surfaces $V(F)$ and $V(G^{\prime})$ intersect transversally at $10$ lines and intersect at the two lines of fixed points with multiplicity $3$. This shows that $Y=X/(\sigma)\to S(\sigma)=Q$ is the blowing down of ten $(-1)$ curves on the Coble surface $Y$.
@@
Here is an example of a bitangent involution. Let $W$ be a general web of quadrics in $\mathbb{P}^{3}$ and let $\mathcal{D}(W)\subset W$ be the discriminant quartic of singular quadrics, classically known as a Cayley quartic symmetroid. A Reye line is a line in $\mathbb{P}^{3}$, which is contained in a pencil from $W$. It is known that the set of Reye lines forms an irreducible congruence ${\text{Rey}}(W)$ of bidegree $(7,3)$, a Reye congruence, and each Reye line is bitangent to the Steinerian surface $X$ of $W$, the set of singular points of quadrics from $W$. The surface $X$ admits a fixed-point-free involution $\sigma$ such that $S(\sigma)={\text{Rey}}(W)$. The involution is defined by $x\mapsto\cap_{Q\in W}P_{x}(Q)$, where $P_{x}(Q)$ is the first polar of $Q$ with pole at $x$.
One can also consider the surface of bitangents ${\mathrm{Bit}}(\mathcal{D}(W))$ of the quartic symmentroid $\mathcal{D}(W)$. It follows from [5, Theorem 7.4.7] that it is an irreducible surface with the normalization isomorphic to the Reye congruence ${\text{Rey}}(W)$. So, ${\text{Rey}}(W)$ admits two birational models as congruences of bidegree $(7,3)$ and $(12,28)$.
Note that the degree of the focal surface $\Phi$ of ${\mathrm{Bit}}(X)$ is equal $22m-2+2g-$ 2, where $g$ is the sectional genus of ${\mathrm{Bit}}(X)$. Since $m=12$, the degree $\geq 22$. However, $\Phi$ could be reducible, and $X$ be its irreducible component. So, the existence of a bitangent involution on $X$ implies that the focal surface of ${\mathrm{Bit}}(X)$ is reducible.
Table (2) gives examples of families of quartic surfaces which admit bitangent involutions with one irreducible component of order $2$.
@@
Of course, any birational involution of a quartic surface $X$ defines an irreducible congruence of lines. If its order is equal to one, it lifts to a birational involution of $\mathbb{P}^{3}$. Given a general point $P$ in $\mathbb{P}^{3}$, there is a unique line passing through $P$ and spanned by an orbit of two points $x,\sigma(x)$ on $X$. We define $T(P)$ to be the fourth point on this line such that the pairs $(x,\sigma(x))$ and $(P,T(P))$ are harmonically conjugate.
@@ 5. Quartic surfaces: $p=2$
Let us see what could be wrong if we drop the assumption that the characteristic is equal to zero.
@@
(supplied by G. Kemper) Take a univariate polynomial $f=a_{n}x^{n}+\cdots+a_{1}x+a_{0}$ whose coefficients are indeterminates over the field $\mathbb{F}_{2}$. With $y_{1},\ldots,y_{n}$ further indeterminates, the polynomial $\prod_{i<j}(y_{i}-y_{j})$ is invariant under the symmetric group $\mathfrak{S}_{n}$ (because the characteristic is two). Since the Galois group of $f$ permutes the roots $r_{i}$, this means that $\prod_{i<j}(r_{i}-r_{j})$ lies in $\mathbb{F}_{2}[a_{1},\ldots,a_{n}]$. (In fact, just write $\prod_{i<j}(y_{i}-y_{j})$ in terms of the elementary symmetric polynomials.) So, the discriminant of $f$ is a square. ∎
@@
The order $m$ of ${\mathrm{Bit}}(X)$ satisfies
$m\leq\frac{1}{4}d(d-1)(d-2)(d-3).$
@@
In characteristic zero, we know that any surface with at most rational double points as singularities is a general projection surface. In the next section, we show that it is not true anymore in characteristic $2$.
@@
It is invariant with respect to the involution
@@
It has one singular point $x_{0}=[1,1,1,1]$ of type $A_{2}$ and four other singular points $[1,\zeta,\zeta^{3},\zeta^{2}]$, where $\zeta^{5}=1,\zeta\neq 1$. The fixed locus $X^{\sigma}$ of $\sigma$ is the line $\ell=V(x_{0}+x_{2},x_{1}+x_{3})$.
For any point $x=[a,b,c,d]\in X$ not lying on $\ell$, the line $\ell_{x}=\langle x,\sigma(x)\rangle$ contains one extra point $[a+c,b+d,a+c,b+d]\in\ell$. It is tangent to $X$ at this point. So, $\sigma$ is an involution of Type (ii).

### v3

@@ 3. Quartic surfaces: $p\neq 2$
In the special case $d=4$, we expect that, for a smooth or nodal quartic surface in characteristic zero, the bidegree $(m,n)$ of ${\mathrm{Bit}}(X)$ is equal to $(12,28)$. Let $S$ be an irreducible component of ${\mathrm{Bit}}(X)$ and let $q_{S}:Z_{S}\to S$ be the restriction of the tautological line bundle $q:Z\to G_{1}(\mathbb{P}^{3})$ to the congruence $S$. The pre-image $\tilde{X}$ of $X$ under the projection $p_{S}:Z_{S}\to\mathbb{P}^{3}$ is a double cover of $S$ defined by the projection $q_{S}$:
@@
Thus, any irreducible component $S$ of ${\mathrm{Bit}}(X)$ with non-zero order and non-zero class defines a birational involution of the surface $\tilde{X}$ with the quotient isomorphic to $S$. In the case when the projection $p_{S}$ restricts to a birational isomorphism $\tilde{X}\to X$, we obtain a birational involution $\sigma_{S}$ of $X$. The set of fixed points of $\sigma_{S}$ is equal to the pre-image of the locus of points on $X$ such that there exists a line intersecting $X$ at this point with multiplicity $4$.
@@
It follows from Kummer’s classification of congruences of order $2$ without fundamental curves that the moduli space of quartic surfaces with $16\geq\mu\geq 11$ nodes contains an irreducible component such that a general surface from this component has reducible congruence ${\mathrm{Bit}}(X)$. These irreducible components are realized as the focal surfaces of the congruences of order 2 without fundamental curves (for example, as we will explain below, the first case in the following Table 2 is a quartic del Pezzo surface). The number of irreducible components and their bidegree are given in Table 2 below (see [15], [23, Art. 348]).
@@
We have already mentioned in the Introduction that a Kummer surface contains 16 planes cut out the surface along a smooth conic. Recall that the Kummer quartic surface is the quotient of the Jacobian of a smooth curve of genus two by the inversion automorphism. The sixteen double conics (tropes) are the images of the theta divisor and its translations by sixteen 2-torsion points. This is a classical, well-known fact that can be found in any treatment of Kummer surfaces (see, for example, [3, 10.3], [9]). Let us explain how to see the six irreducible components of bidegree $(2,2)$. Recall that $X$ admits a smooth model $\tilde{X}$ as a surface of degree $8$ in $\mathbb{P}^{5}$ (see [3, 10.3.3]). The surface $\tilde{X}$ is a complete intersection of three quadrics in $\mathbb{P}^{5}$ which is a $K3$ surface. One can choose projective coordinates such that $\tilde{X}$ is given by the following equations:
@@
$[z_{1},z_{2},z_{3},z_{4},z_{5},z_{6}]\to[\pm z_{1},\pm z_{2},\pm z_{3},\pm z_{4},\pm z_{5},\pm z_{6}]$
induce automorphisms of $\tilde{X}$ which generate a $2$-elementary group of order $2^{5}$. The group of such automorphisms contains a subgroup of index $2$ that consists of the identity and $15$ projective involutions that restrict to symplectic automorphisms of $\tilde{X}$. [^1] This means that the involution leaves invariant a non-zero regular $2$-form on the surface, or, equivalently, the set of its fixed points is non-empty and finite. They are the involutions that change 2 or 4 signs at the coordinates and non-symplectic involutions are the ones that change 1, 3 or 5 signs. The quotient surface of $\tilde{X}$ by a symplectic automorphism is a K3 surface with rational double points. The remaining sixteen involutions are decomposed into two sets. One set consists of ten involutions that change exactly three signs at the coordinates. They are fixed-point-free with the orbit space isomorphic to an Enriques surface. Another set consists of six involutions $g_{i}$ that change only one coordinate with the orbit space isomorphic to a del Pezzo surface of degree $4$. These are the involutions we are interested in.
The first quadric $V(\sum z_{i}^{2})$ can be identified with the Grassmannian quadric $G_{1}(\mathbb{P}^{3})$ written in Klein coordinates $z_{i}$ corresponding to six apolar line complexes (see, for example, [3, 10.2]). It is known that any automorphism of $G_{1}(\mathbb{P}^{3})$ comes from a projective collineation or correlation. The involutions $g_{i}$ preserve the apolar line complexes $V(z_{i})$, and hence, come from a correlation. They transform points to planes, and therefore, act on $G_{1}(\mathbb{P}^{3})$ by transforming the plane of lines through a point (an $\alpha$-plane) to the plane of lines contained in a given plane (a $\beta$-plane).
@@
First, we recall the relation between $X,\tilde{X}$ with the quadratic line complex $Y=V(\sum a_{i}z_{i}^{2})\cap G_{1}(\mathbb{P}^{3})$ (see, e.g., [3, §10.3.3]). For a point $x\in\mathbb{P}^{3}$, let $\Omega(x)\subset G_{1}(\mathbb{P}^{3})$ be the $\alpha$-plane consisting of lines passing through $x$, and for a plane $\Pi\subset\mathbb{P}^{3}$, denote by $\Omega(\Pi)$ the $\beta$-plane consisting of lines lying on $\Pi$. The Kummer quartic surface $X$ is the set of $x\in\mathbb{P}^{3}$ such that $\Omega(x)\cap Y$ is a singular conic in $\Omega(x)$. The surface $\tilde{X}$ sits in $G_{1}(\mathbb{P}^{3})$ as a singular surface of the quadratic line complex, that is, it parameterizes the lines that are singular points of the intersection of $Y$ with $\alpha$-planes. We can identify each point $x\in X$ with the $\alpha$-plane $\Omega(x)$. An automorphism $g_{i}$ acts on $X$ by sending $\Omega(x)$ to $\Omega(\Pi)$. Let $x\in X$ be a general point and let $\Pi=\mathbb{T}_{x}(X)$ be the embedded tangent space of $X$ at $x$. As we explained above, $g_{i}$ interchanges the two families $\{\Omega(x)\}_{x\in\mathbb{P}^{3}}$, $\{\Omega(\Pi)\}_{\Pi\subset\mathbb{P}^{3}}$ of planes, and hence,
@@
$\emptyset\neq\Omega(x)\cap H_{i}=g_{i}(\Omega(x)\cap H_{i})=\Omega(\Pi^{\prime})\cap H_{i},$
$\emptyset\neq\Omega(\Pi)\cap H_{i}=g_{i}(\Omega(\Pi)\cap H_{i})=\Omega(x^{\prime})\cap H_{i}.$
@@
Also, note that the fixed locus of the involution $g_{i}$ is a canonical curve $C$ of genus $5$ given as the intersection of three diagonal quadrics. The curve $C$ has a special group of automorphisms isomorphic to the $2$-elementary group $2^{4}$. It is classically known as a Humbert curve of genus $5$. The image $B$ of $C$ in the quartic del Pezzo surface $\mathcal{D}$ belongs to $|-2K_{\mathcal{D}}|$. The curve $B$ is the branch curve of the double cover $\tilde{X}\to\mathcal{D}$. It has the distinguished property that any of the 16 lines on $\mathcal{D}$ splits into a pair of lines on $\tilde{X}$.
@@ 4. Birational involutions of a quartic surface
Let $\sigma$ be a birational involution of a quartic surface $X$. The closure of lines spanned by orbits of $\sigma$ is an irreducible congruence $S(\sigma)$ of lines. Fix a general point $P\in\mathbb{P}^{3}$, the order of $S(\sigma)$ is equal to the cardinality of the set $\{x\in\mathbb{P}^{3}:P\in\langle x,\sigma(x)\rangle\}$.
@@
where $P=[a,b,c,d],x=[x_{0},x_{1},x_{2},x_{3}],T(x)=[y_{0},y_{1},y_{2},y_{3}]$. It is expected to be a curve of degree $12k+$ 1, where $k$ is the algebraic degree of $T$ [10, p. 175]. However, it also includes the locus of fundamental points of $T$, as well as the closure of the locus of fixed points of $T$. The class of $S(\sigma)$ is expected to be equal to the degree of $\Pi\cap T(\Pi)\cap X$ for a general plane $\Pi$, which is $4k$.
@@
Let $S(\sigma)\subset G_{1}(\mathbb{P}^{3})$ be the congruence of lines defined as the closure of the image of the rational map $\phi_{\sigma}$.
@@
Here, we give an example of an involution of Type (ii). Let $\sigma$ be the restriction of a projective involution $T$ in $\mathbb{P}^{3}$. Suppose that $p\neq 2$ and the fixed locus of $T$ is the union of two skew lines $\ell_{1},\ell_{2}$. It is clear that the lines $\ell_{x}$ are invariant lines of $T$. Hence, each $\ell_{x}$ has two fixed points on it, one on $\ell_{1}$ and one on $\ell_{2}$. This shows that the congruence $S(\sigma)$ is of bidegree $(1,1)$ isomorphic to a smooth quadric. It is known that the set $X^{\sigma}$ of fixed points of an involution $\sigma$ of a K3 surface consists of $8$ isolated fixed points if $\sigma$ is a symplectic involution or it does not have isolated fixed points if $\sigma$ is an anti-symplectic involution. This implies that $T$ intersects $X$ either at $\leq 8$ points (one can show that the equality holds since none of the lines can pass through a singular point, however, we will not need this fact) or $T$ is contained in $X$. In the former case, each line $\ell_{x}$ has two fixed points, one on $\ell_{1}$ and another on $\ell_{2}$. Thus, the degree of $\phi_{\sigma}$ is equal to one.
In the second case, assume $X$ is smooth and $X^{\sigma}$ consists of two lines $\ell_{1}$ and $\ell_{2}$. Then, $Y=X/(\sigma)$ is also smooth. The cover $X\to Y=X/(\sigma)$ is ramified over $\ell_{1}+\ell_{2}$, and its branch curve is the union of two disjoint smooth rational curves $C_{1},C_{2}$ with self-intersection $-4$. By the adjunction formula, $|-K_{Y}|=\emptyset,|-2K_{Y}|=\{C_{1}+C_{2}\}$. Thus, $|-K_{Y}|=\emptyset$ and $|-2K_{Y}|\neq\emptyset$, i.e., $Y$ is a Coble surface obtained by blowing up 10 points on a smooth quadric $Q$ (see [5, §9.1]), the eight intersection points of two nodal quartic curves $\bar{C}_{1},\bar{C_{2}}$ of bidegree $(2,2)$ and the two nodes. The pre-images of the ten exceptional curves on $Y$ in $X$ are ten invariant lines lying in $X$ that contain infinitely many orbits of $\sigma$.
@@
[…]
@@
[…]
A Reye line is a line in $\mathbb{P}^{3}$ which is contained in a pencil from $W$. It is known that the set of Reye lines forms an irreducible congruence ${\text{Rey}}(W)$ of bidegree $(7,3)$ [2, Prop. 3.4.2], a Reye congruence, and each Reye line is bitangent to the Steinerian surface $X$ of $W$.
The surface $X$ admits a fixed-point-free involution $\sigma$ such that $S(\sigma)={\text{Rey}}(W)$. The involution is defined by $x\mapsto\cap_{Q\in W}P_{x}(Q)$, where $P_{x}(Q)$ is the first polar of $Q$ with pole at $x$.
[…]
A point of a congruence of lines is called a fundamental point (in classical terminology, a singular point) if the set of rays passing through this point contains a one-dimensional component. In our case of the congruence ${\mathrm{Bit}}(X)$ a fundamental point must be a singular point of $X$. For example, if $X$ is a Kummer quartic, the rays through its singular point sweep a trope. It is classically known that all rays of a congruence of lines of bidegree $(m,n)$ with only isolated fundamental points are tangent to the focal surface $\Phi$ of the congruence of degree $22m+2g-$ 2, where $g$ is the genus of a general hyperplane section of the congruence (see [23], [11], or [4]). Since $m=12$, the degree is $\geq 22$. However, $\Phi$ could be reducible, and $X$ is its irreducible component. So, the existence of a bitangent involution on $X$ implies that the focal surface of ${\mathrm{Bit}}(X)$ is reducible.
Table 2 gives examples of families of quartic surfaces which admit bitangent involutions with one irreducible component of order $2$.
@@
As we explained, any birational involution $\sigma$ of a quartic surface $X$ defines an irreducible congruence of lines $S(\sigma)$. If the order of $S(\sigma)$ is equal to one, the involution $\sigma$ lifts to a birational involution $T$ of $\mathbb{P}^{3}$. Indeed, given a general point $P$ in $\mathbb{P}^{3}$, there is a unique line passing through $P$ and spanned by an orbit of two points $x,\sigma(x)$ on $X$. We define $T(P)$ to be the fourth point on this line such that the pairs $(x,\sigma(x))$ and $(P,T(P))$ are harmonically conjugate.
@@ 5. Quartic surfaces: $p=2$
[…]
First of all, ${\mathrm{Bit}}(X)$ may not be a surface. For example, let $X$ be a quartic surface in characteristic two with the equation:
$F=x_{0}^{4}+x_{0}^{2}(x_{1}^{2}+A_{2}(x_{2},x_{3}))+x_{1}^{4}+B_{4}(x_{2},x_{3})=0,$
where $A_{2},B_{4}$ are homogeneous polynomials in $x_{2},x_{3}$ of degree 2 and 4, respectively. Let $x=[a,b,0,0]\in V(x_{2},x_{3})$. Substituting in $F$ the parametric equation of a line passing through $x$, we find that the projection map with the center at $x$ is inseparable. Thus, ${\mathrm{Bit}}(X)$ contains a line of $\alpha$-planes.
A point $x$ in $\mathbb{P}^{3}$ is called an inseparable projection center of a normal surface $X$ if the projection map with the center at $x$ is inseparable.
Proposition 5.1.
The set of inseparable projection centers of a normal quartic surface in characteristic $2$ is a finite set or consists of a line. Any inseparable projection center contained in $X$ is a singular point of $X$.
Proof.
Since the set of points $x\in X$ such that the line $\langle q,x\rangle$ is tangent at $x$ is equal to the intersection $X\cap P_{q}(X)$ [3, Theorem 1.1.5], a point $q=[a_{0},a_{1},a_{2},a_{3}]$ is an inseparable center of a quartic surface $X=V(F)$ if and only if the polar $P_{q}(X)=V(\sum a_{i}\frac{\partial F}{\partial x_{i}})=\mathbb{P}^{3}$. Suppose the set of inseparable projection centers contains curve of degree $\geq 2$ (not necessary irreducible). Then we can choose three non-collinear points on it. Choose projective coordinates such that these points are $[1,0,0,0],[0,1,0,0]$ and $[0,0,1,0]$. Then, the partials $\frac{\partial F}{\partial x_{i}}=0$ for $i=0,1,2$. This means that $F$ contains $x_{0},x_{1},x_{2}$ in even power. But, then, it must contain $x_{3}$ in even power. Hence, $F$ is a square.
Suppose $q\in X$ is an inseparable projection center. Then, we may choose projective coordinates such that $q=[0,0,0,1]$ and the equation of $X$ can be written in form $\sum A_{4-k}(x,y,z)w^{k}=0$, where $A_{i}$ is a homogeneous polynomial in $x,y,z$ of degree $i$. Since $q\in X$, $A_{0}=0$ and since $P_{q}(X)=0$, $A_{1}=0$. Thus, $q$ is a singular point. ∎
Corollary 5.2.
Suppose $X$ is nonsingular, then each irreducible componen of ${\mathrm{Bit}}(X)$ is a surface.
@@
[…]
@@
Suppose ${\mathrm{Bit}}(X)$ does not contain $\alpha$-planes. Then, the order $m$ of ${\mathrm{Bit}}(X)$ satisfies
$1\leq m\leq\frac{1}{4}d(d-1)(d-2)(d-3).$
Proof.
The assumption implies that the order of ${\mathrm{Bit}}(X)$ is equal to the number of bitangents dropped from a general point in $\mathbb{P}^{3}$. In the proof of Theorem 2.2, by our assumption, the discriminant of the binary form $g_{d-2}$ is not zero, hence $m\geq 1$. Now apply Proposition 5.3 and obtain the other inequality. ∎
@@
By Corollary 2.5, any surface in characteristic zero with ordinary double points as singularities is a general projection surface. In the next section, we show that it is not true anymore in characteristic $2$.
In the following example of an involution $\sigma$ of Type (ii) of a quartic surface in characteristic $2$ all rays from the congruence of lines $S(\sigma)$ are tangent to the surface.
@@
The surface is invariant with respect to the involution
@@
The set of singular points of $X$ consists of one rational double point $x_{0}=[1,1,1,1]$ of type $A_{2}$ and four ordinary double points $[1,\zeta,\zeta^{3},\zeta^{2}]$, where $\zeta^{5}=1,\zeta\neq 1$. The fixed locus $X^{\sigma}$ of $\sigma$ is the line $\ell=V(x_{0}+x_{2},x_{1}+x_{3})$.
For any point $x=[a,b,c,d]\in X$ not lying on $\ell$, the line $\ell_{x}=\langle x,\sigma(x)\rangle$ contains one extra point $[a+c,b+d,a+c,b+d]\in\ell$. The line $\ell_{x}$ is tangent to $X$ at this point. So, $\sigma$ is an involution of Type (ii).